Solve 6^3n =43^(5n-4).
Thanks

Answer by fractalier(2101)

Okay we are going to take the log, apply the power rules and solve...we get
6^3n = 43^(5n-4)
log 6^3n = log 43^(5n-4)
3n*log 6 = (5n - 4)*log 43
3n*log 6 = 5n*log 43 - 4*log 43
3n*log 6 - 5n*log 43 = 4*log 43
n(3*log 6 - 5*log 43) = 4*log 43
n = 4*log 43 / (3*log 6 - 5*log 43)
now use your calculator for a numerical answer...